OpenStudy (anonymous):

how do you find the domain and range of f(x) when f(x) is 3x-5? and when it's x^2-5?

OpenStudy (anonymous):

f(x)=y Just switch them then replace x and y.

OpenStudy (anonymous):

there is no y in the function?

OpenStudy (anonymous):

f(x)=y.

OpenStudy (anonymous):

When f(x)=3x-5. The doman is x=5/3. When f(x)=x^2-5. The domain is x=squarerootof5.

OpenStudy (anonymous):

When f(x)=x^2-5. The range is y>5

OpenStudy (anonymous):

When f(x)=3x-5. The range is all real numbers.